OpenStudy (anonymous):

identify the y-intercept. y=-3x+2/5 (thats 2 over 5) HELP HELP HELP.

OpenStudy (anonymous):

make x= 0 and u get y= 2/5

OpenStudy (anonymous):

okay, could you maybe explain to me how to do that?

OpenStudy (anonymous):

y= -3*(0) +2/5 y=2/5

OpenStudy (anonymous):

EXAMPLE NOT YOUR MATH PROBLEM SLOPE INTERCEPT FORM Y = MX +B M = SLOPE B = Y INTERCEPT

OpenStudy (anonymous):

wait, whats mx+b?

OpenStudy (anonymous):

EXAMPLE NOT YOUR MATH PROBLEM (I HOPE YOU KNOW THIS) SLOPE INTERCEPT FORM Y = MX + B Y = Y VALUE X = X VALUE M = SLOPE B = Y INTERCEPT

OpenStudy (anonymous):

ohhhhhhh, your awesome. (:

OpenStudy (anonymous):

one last question, how do i put them all together? *test question* y= 2/23 (two over twenty three)x-20